A planet orbiting the sun 42,000 miles from the center of the sun. It makes 1 revolution in 3 days. Find the speed of the planet in miles per hour.
Answer by nerdybill(7384)

A planet orbiting the sun 42,000 miles from the center of the sun. It makes 1 revolution in 3 days. Find the speed of the planet in miles per hour.
.
circumference is (pi)d
where
d is diameter
.
circumference = (3.14)(2)42000 = 263760 miles
.
3 days = 3(24) = 72 hours
.
Speed = 263760/72 = 3663.33 mph